OpenStudy (anonymous):

In a recent year, the number of $1 bills in circulation in the U.S. was about 7 billion. If the number of $10 bills was 20% of the number of $1 bills, about how many $10 bills were in circulation? a. 1.4 billion c. .5 billion b. 5.6 billion d. 10 billion

OpenStudy (jdoe0001):

there were 7billion $1 dollar bills and there were 20% of THAT MUCH $10 bills, how many $10 bills were there? well, 20% of of THAT MUCH will be 20% of 7billion \(\bf \textit{n% of m}=\cfrac{n}{100}\times m\implies 20\%\textit{ of }7billions\implies \cfrac{20}{100}\times 7billions\)

OpenStudy (anonymous):

20% of 7 billion (the number of $1 bills) = the number of $10 bills in circulation. We can write: .2*(7billion)=1.4 billion (this is the number of $10 bills in circulation)